### v1

# Vevo2: Bridging Controllable Speech and Singing Voice Generation via Unified Prosody Learning

Abstract
Controllable human voice generation, particularly for expressive domains like singing, remains a significant challenge. This paper introduces Vevo2, a unified framework for controllable speech and singing voice generation. To tackle issues like the scarcity of annotated singing data and to enable flexible controllability, Vevo2 introduces two audio tokenizers: (1) a music-notation-free prosody tokenizer that captures prosody and melody from speech, singing, and even instrumental sounds, and (2) a low-frame-rate (12.5 Hz) content-style tokenizer that encodes linguistic content, prosody, and style for both speech and singing, while enabling timbre disentanglement. Vevo2 consists of an auto-regressive (AR) content-style modeling stage, which aims to enable controllability over text, prosody, and style, as well as a flow-matching acoustic modeling stage that allows for timbre control. Particularly, during pre-training of the AR model, we propose both explicit and implicit prosody learning strategies to bridge speech and singing voice. Moreover, to further enhance the AR model’s ability to follow text and prosody, we design a multi-objective post-training task that integrates both intelligibility and prosody similarity alignment. Experimental results show that the unified modeling in Vevo2 brings mutual benefits to both speech and singing voice generation. Additionally, Vevo2’s effectiveness across a wide range of synthesis, conversion, and editing tasks for both speech and singing further demonstrates its strong generalization ability and versatility. Audio samples are are available at https://versasinger.github.io/. We will release code and model checkpoints at Amphion [^1] https://github.com/open-mmlab/Amphion Zhang et al. (2024a).
1 Introduction
Controllable generation of human voice is an important research topic in the field of audio generation Tan (2023); Xie et al. (2024). In recent years, remarkable progress has been achieved in speech generation, particularly in zero-shot text-to-speech (TTS) Ju et al. (2024); Anastassiou et al. (2024a); Wang et al. (2025a); Du et al. (2024a), largely propelled by the availability of large-scale training corpora. However, for more challenging domains such as singing voice—which can be regarded as a highly expressive form of speech with strictly constrained prosody (e.g., melodySundberg and Rossing (1990)—achieving both high generation quality and controllability remains challenging Bai et al. (2024); Yuan et al. (2025).
This paper focuses on unifying controllable speech and singing voice generation. Our motivation is based on the hypothesis that learning for speech and singing voice can be mutually beneficial through unified modeling. Specifically, the abundance of speech data can supplement the typically smaller singing voice datasets He et al. (2024); Zhang et al. (2024b), potentially improving the overall quality of generated singing voice Yuan et al. (2025). Conversely, the inherent expressiveness of singing voice, when incorporated into a unified learning framework, can enhance expressive speech generation Gu et al. (2024) and strengthen prosody-following capabilities, which is especially advantageous for tasks such as speech editing (Section 4.3).
However, building such a unified system poses significant challenges. First, corpora for singing voice generation have traditionally relied heavily on expert annotations Umbert et al. (2015); Liu et al. (2022); Zhang et al. (2024c); Zhao et al. (2025), such as detailed music notation with precise alignment between lyrics and musical notes Zhang et al. (2022a); Dai et al. (2023); Zhang et al. (2024b). These datasets are scarce, which severely limits scalability. Moreover, this unique data organization is not well-suited for unified modeling of speech and singing voice. Second, how can we achieve precise controllability within this unified system, enabling independent manipulation of text (or lyrics), prosody (or melody), style (such as accent, emotion, or singing techniques Wilkins et al. (2018); Zhang et al. (2024b)), and timbre (i.e., speaker identity)?
To address these challenges, this study first proposes two distinct audio tokenizers: (1) Prosody Tokenizer: A single-codebook VQ-VAE tokenizer van den Oord et al. (2017); Kumar et al. (2023) operating at 6.25 Hz (56.25 bps), trained to reconstruct the chromagram Müller (2015) of raw audio. This tokenizer unifies the modeling of both melody in singing voices and prosody in speech, and can directly extract tokens from the chromagram of raw audio, eliminating the need for expert-annotated music notation and thus greatly improving scalability. Besides, due to the chromagram’s strong generalization ability across various types of audio data Müller (2015); Copet et al. (2023), this tokenizer can even capture the melody of non-human sounds, such as instrumental music (Section 4.3), thereby enhancing the flexibility of melody control in downstream controllable music generation tasks Bai et al. (2024); Yuan et al. (2025). (2) Content-Style Tokenizer: It employs a similar architecture to the prosody tokenizer, but operates at 12.5 Hz (175 bps) and is trained to reconstruct both chromagram features and Whisper’s hidden features Radford et al. (2023). It effectively captures linguistic content, melody, and style in both speech and singing voice, while also achieving robust timbre disentanglement (Table 3). Its role is similar to the “semantic tokens” in speech generation Borsos et al. (2023); Wang et al. (2025a); Du et al. (2024a); Zhang et al. (2025a), but it applies to both speech and singing voice, and operates at a lower frame rate of 12.5 Hz.
Building upon these proposed tokenizers, we introduce Vevo2, a versatile and controllable voice generation framework designed to unify speech and singing voice. Vevo2 adopts a two-stage architecture, similar to classic zero-shot speech generation systems Anastassiou et al. (2024a); Du et al. (2024a); Zhang et al. (2025a) (Figure 1). It includes an Auto-Regressive (AR) content-style modeling stage, which can be prompted by a style reference, and a Flow-Matching (FM) acoustic modeling stage, which can be prompted by a timbre reference. The uniqueness of Vevo2 lies in its AR stage: (1) Pre-training: To enable unified prosody learning for speech and singing voice, we design two different training strategies: Explicit Prosody Learning (EPL), where prosody tokens are provided as explicit input, allowing the model to predict content-style tokens from both text and prosody tokens; and Implicit Prosody Learning (IPL), where prosody tokens are not used as input, and content-style tokens are predicted solely from text tokens. By randomly applying both EPL and IPL to all voice data, the model learns prosody in a more unified manner, effectively bridging the gap between speech and singing characteristics. (2) Post-training: To further improve the text- and prosody-following abilities of the pre-trained model, we introduce a multi-objective post-training task. We propose a composite reward that incorporates both intelligibility and prosody similarity, and extend the GRPO algorithm Shao et al. (2024) for its optimization. Post-training significantly enhances the model’s controllability with respect to text and melody, and also improves its generalization to out-of-pretraining-distribution data, such as instrumental sounds.
The contributions of this paper are summarized as follows:
  • We propose two audio tokenizers: (1) a unified prosody tokenizer capable of encoding prosody and melody from speech, singing voice, and even instruments sounds, without reliance on music notation; and (2) a low-frame-rate (12.5 Hz) content-style tokenizer that models linguistic content, prosody, and style for both speech and singing voice while achieving timbre disentanglement.
  • We introduce Vevo2, a unified and controllable framework for both speech and singing voice generation. Its effectiveness across diverse synthesis, conversion, and editing tasks demonstrates strong generalization and versatility. Moreover, its superior melody-following capability enables unique music generation applications, such as humming-to-singing and instrument-to-singing.
  • We propose a multi-objective post-training task for controllable human voice generation. By optimizing both intelligibility and prosody similarity rewards with a GRPO-based algorithm, Vevo2 achieves greater stability and improved overall performance across a variety of tasks.
2 Related Work
Controllable Voice Generation Controllable speech and singing voice generation has consistently remained a significant research focus. (1) In the speech domain, existing works explore control over multiple attributes: linguistic content, which is fundamental to TTS Tan (2023); Xie et al. (2024) and speech editing tasks Le et al. (2023); Peng et al. (2024); Wang et al. (2025b); Chen et al. (2025); timbre, which is central to VC tasks Sun et al. (2016); Qian et al. (2019); Ju et al. (2024); Anastassiou et al. (2024b); Zhang et al. (2025a); Liu (2024); and style, which encompasses accent Zhao et al. (2021); Xue et al. (2024); Zhang et al. (2025a), emotion Zhou et al. (2022); Qi et al. (2024); Zhang et al. (2025a), whisper Meenakshi and Ghosh (2018), and a broader natural language description Lyth and King (2024); Ji et al. (2024); Diwan et al. (2025). (2) For singing voice, beyond controlling lyrics, timbre, and style—as explored in SVS Liu et al. (2022); Zhang et al. (2022b, 2024c, 2024d); Guo et al. (2025), SVC Huang et al. (2023); Ning et al. (2023); Sha et al. (2024); Zhang et al. (2024e), and singing lyric editing tasks Zhang et al. (2022c); Hong et al. (2023); Yang et al. (2025)melody control holds particular significance due to the inherent musicality of singing voices Copet et al. (2023); Bai et al. (2024); Yuan et al. (2025). (3) From a unified modeling perspective, previous research has attempted to integrate various speech generation tasks such as TTS and VC Zhang et al. (2025a); Du et al. (2024a); Wang et al. (2025c), or different singing voice generation tasks like SVS and SVC Hong et al. (2023); Dai et al. (2025). However, models capable of unifying both speech and singing voice domains remain scarce Yang et al. (2024a); Lei et al. (2023). Moreover, the potential benefits of joint speech and singing pre-training, along with the versatile control mechanisms for multiple voice attributes within a unified system, remain largely unexplored.
Speech Tokenizer Recent research has emphasized the disentangled speech representation learning, aiming to separately model the factorized speech attributes to reduce learning complexity Jiang et al. (2023); Ju et al. (2024) and achieve more precise controllability Polyak et al. (2021); Ju et al. (2024); Zhang et al. (2025a). Common speech tokenizers include linguistic content tokenizers Hsu et al. (2021); Zhang et al. (2025a); Baade et al. (2025), content-style tokenizers Borsos et al. (2023); Anastassiou et al. (2024a); Wang et al. (2025a); Du et al. (2024a); Zhang et al. (2025a), and prosody tokenizers Jiang et al. (2023); Ju et al. (2024). However, in pursuing a unified controllable model for both speech and singing generation, existing speech tokenizers exhibit several limitations: As these tokenizers predominantly focus on the speech domain, their effectiveness in handling singing voice data remains unexplored. The design of a unified prosody tokenizer capable of handling both modalities, along with enhancing the melody encoding capabilities of content-style tokenizers Zhang et al. (2024e), remains an important issue for future exploration.
Post-training for Audio Generation Alignment via post-training has demonstrated its effectiveness in the generation of text Ouyang et al. (2022); Bai et al. (2022), vision Xu et al. (2023); Fu et al. (2024), speech Zhang et al. (2024f); Anastassiou et al. (2024a); Du et al. (2024b), music Cideron et al. (2024), and sound effects Majumder et al. (2024); Liao et al. (2024). In speech generation, existing works have employed preference alignment to enhance multiple aspects of speech, including intelligibility Anastassiou et al. (2024a); Du et al. (2024b); Zhang et al. (2025b), speaker similarity Anastassiou et al. (2024a); Du et al. (2024b), emotion controllability Anastassiou et al. (2024a); Gao et al. (2024), and overall quality Zhang et al. (2024f); Chen et al. (2024); Yao et al. (2025). For developing a unified speech and singing voice generation model, the importance of specific capability elicitation through post-training becomes particularly crucial, given the expanded range of capabilities injected into the pre-trained model. However, the design of appropriate alignment objectives for unified speech and singing voice generation remains an open research problem.
3 Methodology
Task Input Auto-Regressive Stage Flow-Matching Stage
Text to Speech text, reference wav $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$] $\boldsymbol{\bm{cs}}$, $\boldsymbol{\bm{mel}}$, $\boldsymbol{\bm{\hat{cs}}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{mel}}}$]
Singing Voice Synthesis text, midi, reference wav $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$]
Voice Conversion (style-converted)
source wav, reference wav $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$]
Singing Voice Conversion (style-converted)
$\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$]
Speech Editing edited text, raw wav $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$]
Singing Lyric Editing
Humming to Singing text, humming wav, singer wav $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{p}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$]
Instrument to Singing text, instrument wav, singer wav
Emotion Conversion source wav, style wav $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$] $\boldsymbol{\bm{cs}}$, $\boldsymbol{\bm{mel}}$, $\boldsymbol{\bm{\hat{cs}}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{mel}}}$]
Accent Conversion
Whisper-to-Normal Conversion
Singing Style Conversion $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$]
Voice Conversion (style-preserved)
source wav, reference wav - $\boldsymbol{\bm{cs}}$, $\boldsymbol{\bm{mel}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{mel}}}$]
Singing Voice Conversion (style-preserved)
    We denote the text tokens, prosody tokens, content-style tokens, and acoustic representations (specifically, Mel spectrogram) as $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{cs}}$, and $\boldsymbol{\bm{mel}}$, respectively. We use blue, red, and green to indicate the source inputs of these representations. $\boldsymbol{\bm{\hat{cs}}}$ and $\boldsymbol{\bm{\hat{mel}}}$ represent the model-predicted outputs. For the singing voice synthesis (SVS) task, we render MIDI notations into instrumental sounds and then extract their prosody tokens (Section 4.3).
Figure 1: Vevo2 inference pipeline for versatile synthesis, conversion, and editing tasks.
3.1 Prosody Tokenizer
Motivation In order to model both the prosody of speech and the melody of singing voice using a unified method, we choose chromagram feature Müller (2015) and design a VQ-VAE tokenizer van den Oord et al. (2017); Kumar et al. (2023) based on it. Some previous studies of speech generation tried to use fundamental frequency (F0) Polyak et al. (2021); Ju et al. (2024) or the low-frequency band of the mel spectrogram Jiang et al. (2023) to build a prosody tokenizer. The reasons why we adopt the chromagram are as follows: (1) Octave-free, thus friendly for unification: There is a significant distribution gap of F0 between speech and singing voice, such as the F0 range of singing voice is much wider Zhang et al. (2023). Chromagram removes octave information between singing voice and speech, making it easier to capture common prosodic patterns across both domains. (2) Notation-free, thus friendly for scaling up: Chromagram effectively represents high-level musical and singing voice attributes such as melody and harmony Müller (2015), and can be automatically extracted from audio without manual transcription or music notation, making it suitable for large-scale data processing. (3) Robust for diverse audio domains: As a statistical distribution of spectral energy, chromagram is less sensitive to noise and pitch detection errors De Cheveigné and Kawahara (2002), resulting in more stable performance. In addition, chromagram provides strong representations for both human voice and non-human sounds, such as instrumental music Müller (2015); Copet et al. (2023). These characteristics enhance the flexibility of melody control in downstream music generation tasks, such as enabling instrument-to-sing applications (Section 4.4).
Architecture The prosody tokenizer consists of five main components: Downsampling Layer, Encoder, Vector Quantization (VQ), Decoder, and Upsampling Layer. Formally, given the codebook $\bm{E}=[\bm{e}_{1},\bm{e}_{2},\dots,\bm{e}_{K}]$ with vocabulary size $K$, and taking chromagram features $\bm{x}$ as input, the reconstructed output $\hat{\bm{x}}$ is obtained after passing through these five modules:
$\begin{split}\bm{z}_{e}(\bm{x})&=\text{Encoder}(\text{Downsample}(\bm{x})),\\ \bm{z}_{q}(\bm{x})&=\bm{e}_{k},~\text{where}~k=\arg\min_{j}\|\bm{z}_{e}(\bm{x})-\bm{e}_{j}\|_{2},\\ \hat{\bm{x}}&=\text{Upsample}(\text{Decoder}(\bm{z}_{q}(\bm{x}))),\end{split}$ (1)
where $\bm{z}_{q}(\bm{x})$ is the quantized representation (i.e., token) of $\bm{z}_{e}(\bm{x})$ after VQ. The loss function consists of the reconstruction loss (whose weight is $\lambda$) and quantization loss (whose weight is $\beta$):
$\mathcal{L}=\lambda\|\bm{x}-\hat{\bm{x}}\|^{2}_{2}+\beta\|\bm{z}_{e}(\bm{x})-\bm{z}_{q}(\bm{x})\|^{2}_{2}.$ (2)
We follow Yu et al. (2022); Kumar et al. (2023); Wang et al. (2025a) and adopt factorized (i.e., a low-dimensional code-lookup space) and L2-normalized strategies for VQ learning. We extract 50 Hz chromagram features from raw audio and apply an 8x downsampling, resulting in a prosody tokenizer operating at 6.25 Hz (56.25 bps).
3.2 Content-Style Tokenizer
Motivation We follow the main ideas from recent works in controllable speech generation Zhang et al. (2025a); Du et al. (2024a), and aim to develop a content-style tokenizer that can encode linguistic content, prosody, and style, while also achieving timbre disentanglement. Specially, to unify both speech and singing voice, beyond encoding linguistic content and style, as in typical speech generation tasks Zhang et al. (2025a); Du et al. (2024a); Anastassiou et al. (2024a); Wang et al. (2025a), we also require the tokenizer to capture melody and fine-grained style information (e.g., F0 and energy details) in singing voice. Motivated by this, when designing our VQ-VAE tokenizer, we not only use hidden features from automatic speech recognition (ASR) models—specifically, from Whisper Radford et al. (2023)—as reconstruction objectives, but also incorporate chromagram features as an additional reconstruction objective. This compensates for the limitations of Whisper features in modeling melody in singing voice Zhang et al. (2024e). Furthermore, to facilitate the downstream content-style modeling stage, we design the content-style tokenizer to operate at a low frame rate (specifically, 12.5 Hz in our study, which is lower than the commonly used 25 or 50 Hz Anastassiou et al. (2024a); Zhang et al. (2025a); Wang et al. (2025a); Du et al. (2024a, b)). This reduces the sequence length for the AR model and thus alleviates its learning burden.
Architecture The architecture of the content-style tokenizer is similar to the proposed prosody tokenizer. The only difference lies in the input and reconstruction objectives incorporate both chromagram features and Whisper features. We extract 50 Hz chromagram features and 50 Hz Whisper features (which is the encoder output of the pre-trained Whisper model Radford et al. (2023)) and apply a 4x downsampling, resulting in a content-style tokenizer operating at 12.5 Hz (175 bps).
3.3 Pre-training
Content-Style Modeling For content-style modeling, the common approach in speech generation works is to learn a mapping from text tokens to content-style tokens (typically using an auto-regressive Borsos et al. (2023); Anastassiou et al. (2024a); Du et al. (2024a); Zhang et al. (2025a) or masked generative model Wang et al. (2025a)). We refer to this as Implicit Prosody Learning (IPL), since the model implicitly supplements prosody information for the mapping through in-context learning. However, for singing voice generation, it is often desirable to explicitly control the prosody (i.e., melody) of the generated output Umbert et al. (2015); Liu et al. (2022); Zhang et al. (2024c); Zhao et al. (2025). To address it, we propose an Explicit Prosody Learning (EPL) strategy, where prosody tokens are provided as explicit input, allowing the model to predict content-style tokens from both text and prosody tokens. Notably, during training, we do not restrict IPL to speech data and EPL to singing voice data. Instead, we treat all samples equally: for each training sample, we randomly choose between IPL and EPL with equal probability. We assume that this design better unifies the training patterns for speech and singing voice.
During pre-training, we initialize the AR transformer with a pre-trained textual large language model (LLM), specifically Qwen2.5-0.5B Yang et al. (2024b). Given a speech or singing voice sample u{u}, we denote the prosody and content-style tokens as $\bm{Q}_{p}({u})$ and $\bm{Q}_{cs}({u})$, respectively. The input sequences to the transformer for EPL and IPL are shown in Equation 3 and Equation 4, where $\bm{I}_{ipl}$ and $\bm{I}_{epl}$ represent the text instruction prefixes for IPL and EPL (see our Appendix for more details), $\bm{T}$ denotes the text of u{u}, and $\langle|\cdot|\rangle$ indicates special tokens. We perform next token prediction only on the final segment $\langle|\text{start-of-cs}|\rangle,\bm{Q}_{cs}({u}),\langle|\text{end-of-cs}|\rangle]$:
$\bm{I}_{ipl},\bm{T},\langle|\text{start-of-cs}|\rangle,\bm{Q}_{cs}({u}),\langle|\text{end-of-cs}|\rangle].$ (3)
$\bm{I}_{epl},\bm{T},\langle|\text{start-of-p}|\rangle,\bm{Q}_{p}({u}),\langle|\text{end-of-p}|\rangle,\langle|\text{start-of-cs}|\rangle,\bm{Q}_{cs}({u}),\langle|\text{end-of-cs}|\rangle].$ (4)
During inference, when the prosodic source is not used, the inference paradigm is identical to classic zero-shot TTS Wang et al. (2023); Anastassiou et al. (2024a); Zhang et al. (2025a); Du et al. (2024a). As illustrated in the TTS task in Figure 1, the process can be represented as $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$], where the style signals in $\boldsymbol{\bm{\hat{cs}}}$ are prompted by $\boldsymbol{\bm{cs}}$ and generated through the in-context learning capabilities of the pre-trained model. Particularly, when both prosodic source and style reference are utilized as inputs, the inference paradigm takes the form $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$]. In this context, for $\boldsymbol{\bm{\hat{cs}}}$, we expect $\boldsymbol{\bm{p}}$ to provide coarse-grained prosody contour signals, while $\boldsymbol{\bm{cs}}$ contributes fine-grained style details (e.g., the vibrato patterns for singing). We will discuss this further through our experiments on style-converted singing voice conversion (Section 4.3).
Acoustic Modeling During acoustic modeling, we follow the common approach adopted in speech generation studies Bai et al. (2024); Du et al. (2024a); Zhang et al. (2025a); Zeng et al. (2025), employing a flow-matching transformer to convert content-style tokens into Mel spectrograms, conditioned on a timbre reference. After obtaining the Mel spectrogram, we use a Vocos-based vocoder Siuzdak (2024) that is pre-trained on both speech and singing voice, to generate the final waveform. We provide the detailed architecture designs in Appendix A.3.
3.4 Post-training
Motivation After pre-training, we observe that the AR model exhibits good versatility across diverse tasks (Section 4.3). Nevertheless, its stability remains suboptimal, particularly in its fundamental capabilities of text- and prosody-following, which can be further elicited and enhanced. Furthermore, in downstream applications, the model may encounter out-of-pretraining data such as instrumental sounds. For instance, in the context of SVS, the model needs to accept prosody tokens extracted from MIDI-rendered sounds for melody control (Figure 1). Motivated by the above, this study further employs post-training to enhance the model’s controllability for text and prosody, as well as to improve its robustness to diverse audio data. We propose a multi-objective post-training task that jointly aligns both intelligibility and prosody similarity (Figure 2). We believe that this joint alignment helps prevent the model from over-optimizing a single objective (as evidenced in Section 4.4).
Figure 2: Multi-objective alignment for both intelligibility and prosody similarity. This figure demonstrates how we utilize instrumental music as prosody prompts during post-training of Vevo2.
Intelligibility Reward To align intelligibility, we adopt a recent zero-shot TTS intelligibility preference corpus, INTP Zhang et al. (2025b), which contains diverse intelligibility preference pairs constructed through comparisons between multiple models as well as manually designed intelligibility perturbations, to train an intelligibility reward model. Specifically, we employ the commonly used Bradley-Terry (BT) approach Bradley and Terry (1952); Ouyang et al. (2022); Bai et al. (2022) for reward model training. Given an intelligibility preference dataset $\mathcal{D}=\{(t,a^{w},a^{l})\}$, where $a^{w}$ and $a^{l}$ represent the positive and negative audio outputs with respect to the target text $t$, we parametrize an intelligibility reward model $r_{\phi}(t,a)$, treat the problem as binary classification, and optimize the negative log-likelihood loss:
$\mathcal{L}_{R}(r_{\phi},\mathcal{D})=-\mathbb{E}_{(t,a^{w},a^{l})\sim\mathcal{D}}\left[\log\sigma\left(r_{\phi}(t,a^{w})-r_{\phi}(a,y^{l})\right)\right],$ (5)
where $r_{\phi}(t,a)$ is initialized from the pre-trained model. We follow the pre-training and use both EPL and IPL for training this reward model. During inference, we can input it a $\bm{T},\bm{Q}_{cs}]$ or $\bm{T},\bm{Q}_{p},\bm{Q}_{cs}]$ sequence to obtain an intelligibility reward. More details are provided in Appendix A.4.1.
Prosody Similarity Reward To achieve prosodic alignment, we establish the following protocol to evaluate the policy model’s prosody-following capability: (1) Suppose we have a singing voice dataset with MIDI annotations. Given a singing voice sample u{u}, we use its MIDI-rendered instrumental music to extract prosody tokens as control conditions, denoted as $\bm{Q}_{p}$; (2) Let the policy model perform a completion conditioned on $\bm{Q}_{p}$ and $\bm{T}$ (where $\bm{T}$ can be any text), and denote the generated content-style tokens as $\hat{\bm{Q}}_{cs}$; (3) Using the decoder of the content-style tokenizer (Section 3.2), we reconstruct the chromagram features $\hat{\bm{x}}$ from $\hat{\bm{Q}}_{cs}$, and compute the cosine similarity between $\hat{\bm{x}}$ and the ground-truth chromagram features $\bm{x}$ of the singing voice u{u}. [^2] Note that in step (3), we could also use the flow-matching model and vocoder to perform inference and extract $\hat{\bm{x}}$ from the predicted waveform, but for training efficiency, we adopt the current approach. In practice, our prosody prompt data includes not only the aforementioned instrumental music data, but also some of the speech and singing voice data used during pre-training. More details are provided in Appendix A.4.2.
Multi-Objective Optimization We apply the GRPO algorithm Shao et al. (2024) to optimize the proposed multi-objective alignment task. Assuming the number of completions per prompt is $K$, for each prompt, after obtaining the intelligibility rewards ($\bm{r}_{int}$) and prosody similarity rewards ($\bm{r}_{pro}$) for all $K$ completions, we apply group normalization to both rewards separately. The sum of these normalized rewards serves as the final outcome supervision:
$A^{(i)}=\frac{r_{int}^{(i)}-\text{mean}(\bm{r}_{int})}{\text{std}(\bm{r}_{int})}+\frac{r_{pro}^{(i)}-\text{mean}(\bm{r}_{pro})}{\text{std}(\bm{r}_{pro})},$ (6)
where $r_{int}^{(i)}$ and $r_{pro}^{(i)}$ denote the intelligibility reward and prosody similarity reward for the $i$-th completion, respectively, and $A^{(i)}$ represents the advantage.
3.5 Inference-Time Control
After obtaining the pre-trained (or post-trained) AR model and FM model, we can perform inference following the pipeline illustrated in Figure 1, by utilizing different combinations of text, prosodic source, style reference, and timbre reference to achieve versatile synthesis, conversion, and editing tasks for both speech and singing voice. In this section, we further explain two additional inference-time control mechanisms of Vevo2: Duration Control and Pitch Region Control.
Duration Control Some Non-AR based speech generation models have often featured the ability to control the total duration of generated speech Wang et al. (2025a); Chen et al. (2025), while AR-based models are generally considered less capable of controlling output duration Borsos et al. (2023); Anastassiou et al. (2024a); Du et al. (2024a); Zhang et al. (2025a); Zhou et al. (2025). Interestingly, however, we discover that Vevo2 can effectively control the duration of generated results through the use of prosody tokens as input. This capability stems from our tokenizers’ fixed frame rates (6.25 Hz and 12.5 Hz), which maintain a consistent 2:1 ratio between content-style and prosody token sequences during training. Consequently, during inference, we can control the duration of the generated output by applying simple linear scaling transformations to the chromagram features of the prosodic source, thereby modifying the length of prosody tokens. Such control mechanism is detailed in Appendix A.5.1.
Pitch Region Control For voice conversion (VC) or singing voice conversion (SVC) tasks, due to potential significant differences in timbre types Wilkins et al. (2018); Sundberg and Rossing (1990) between source and reference speakers, we often need to adjust the pitch region (i.e., apply pitch shift) of the source waveform during inference to achieve higher speaker similarity Huang et al. (2023); Zhang et al. (2024e); Sha et al. (2024). In Vevo2, we discover that pitch region control can also be achieved: by shifting the pitch of the source waveform before prosody token extraction, we can effectively control the pitch region of the generated speech or singing voice. Our experiments demonstrate that this approach is effective both when using only the FM stage and when combining AR and FM stages. Such control mechanism is detailed in Appendix A.5.2.
4 Experiments
Table 1: Results on zero-shot TTS task. (PT?: Whether post-training is applied.)
Model
#Hz
Data
(#hours)
PT?
Expressive Speech Singing Voice
WER SIM
N-
CMOS
$S$-
CMOS
WER SIM
N-
CMOS
$S$-
CMOS
Ground Truth - - - 10.91 0.687 1.47 $_{\scriptscriptstyle\pm\text{0.24}}$ 0.87 $_{\scriptscriptstyle\pm\text{0.19}}$ 12.65 0.658 1.04 $_{\scriptscriptstyle\pm\text{0.18}} -$0.24 $_{\scriptscriptstyle\pm\text{0.15}}$
F5-TTS Chen et al. (2025) - 101K 11.77 0.695 -1.05 $_{\scriptscriptstyle\pm\text{0.18}} -$0.06 $_{\scriptscriptstyle\pm\text{0.17}}$ 16.12 0.597 -1.92 $_{\scriptscriptstyle\pm\text{0.06}} -$1.53 $_{\scriptscriptstyle\pm\text{0.08}}$
MaskGCT Wang et al. (2025a) 50 101K 13.42 0.736 -1.14 $_{\scriptscriptstyle\pm\text{0.21}}$ 0.07 $_{\scriptscriptstyle\pm\text{0.31}}$ 11.71 0.753 -1.83 $_{\scriptscriptstyle\pm\text{0.14}} -$1.14 $_{\scriptscriptstyle\pm\text{0.11}}$
CosyVoice 2 Du et al. (2024a) 25 167K 11.20 0.706 0.10 $_{\scriptscriptstyle\pm\text{0.29}} -$0.16 $_{\scriptscriptstyle\pm\text{0.13}}$ 16.18 0.659 -1.63 $_{\scriptscriptstyle\pm\text{0.17}} -$0.94 $_{\scriptscriptstyle\pm\text{0.13}}$
Vevo2-base
12.5
101K 15.52 0.677 -0.84 $_{\scriptscriptstyle\pm\text{0.15}} -$0.06 $_{\scriptscriptstyle\pm\text{0.27}}$ 19.39 0.658 -1.15 $_{\scriptscriptstyle\pm\text{0.13}} -$0.65 $_{\scriptscriptstyle\pm\text{0.18}}$
7K 37.57 0.611 -1.25 $_{\scriptscriptstyle\pm\text{0.26}} -$0.57 $_{\scriptscriptstyle\pm\text{0.17}}$ 25.38 0.692 -0.15 $_{\scriptscriptstyle\pm\text{0.23}} -$0.23 $_{\scriptscriptstyle\pm\text{0.14}}$
101K, 7K 14.32 0.681 -0.49 $_{\scriptscriptstyle\pm\text{0.24}} -$0.03 $_{\scriptscriptstyle\pm\text{0.19}}$ 15.78 0.708 -0.66 $_{\scriptscriptstyle\pm\text{0.16}} -$0.17 $_{\scriptscriptstyle\pm\text{0.24}}$
Vevo2 12.5 101K, 7K 11.48 0.689 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 7.66 0.725 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$
    #Hz: Frame rate of the content-style tokenizer. Data: Pre-training data, where blue and red denote speech and singing data. The results on regular speech domain are in Appendix C.1. The best and the second best result (excluding ground truth) is shown in bold and by underlined.
4.1 Experimental Setup
Training Data For pre-training, we use Emilia He et al. (2024) as the speech data (101K hours), and follow SingNet’s pipeline Gu et al. (2025) to prepare 7K hours of singing voice data (source-separated from in-the-wild songs from the Internet). For post-training, we utilize INTP Zhang et al. (2025b) (which contains 250K preference speech pairs) to train the intelligibility reward model. For the proposed multi-objective alignment task, we employ three types of prosody prompts: (1) speech data: 20K positive samples randomly selected from INTP, (2) singing voice data: 20K samples from M4Singer Zhang et al. (2022a), and (3) instrumental sounds: rendered from MIDIs in M4Singer, covering 16 instruments.
Evaluation Data We consider various settings to construct the evaluation set: (1) For regular speech, we utilize SeedTTS test sets Anastassiou et al. (2024a), which is widely adopted in the TTS field Wang et al. (2025a); Du et al. (2024b); Chen et al. (2025). (2) For expressive speech, we employ Genshin-Voice Simon3000 (2025), which consists of professional voice acting recordings from the game Genshin Impact. (3) For singing voice, we use GTSinger Zhang et al. (2024b) and SingStyle111 Dai et al. (2023), which encompass a wide range of singing techniques, styles, and timbre types. (4) Additionally, for two unique tasks: humming-to-singing and instrument-to-singing, we utilize humming Liu et al. (2024) and instrumental sound including flute Braga Brum (2018) and violin Dong et al. (2022) for the evaluation. All evaluations are conducted using target text in both English and Chinese languages.
[…]
4.2 Mutual Benefits of Unified Modeling for Speech and Singing Voice
To investigate whether speech and singing voice generation can be mutually beneficial through unified modeling, we examine the impact of different pre-training data distributions on Vevo2. Specifically, we conduct experiments on zero-shot TTS tasks and evaluate the performance using reference prompts from different distributions. For comparison, we select models with varying content-style tokenizer frame rates: MaskGCT Wang et al. (2025a) (50 Hz), CosyVoice 2 Du et al. (2024b) (25 Hz), and F5-TTS Chen et al. (2025) (a single-stage model without content-style tokenizer). The results are presented in Table 1.
The experimental results demonstrate the following findings: (1) Comparing Vevo2 models with different pre-training data, we observe that on expressive speech domain, jointly using speech and singing data for pre-training improves all metrics (WER: 15.52 $\rightarrow$ 14.32, SIM: 0.677 $\rightarrow$ 0.681, N-CMOS: -0.84 $\rightarrow$ -0.49, SS-CMOS: -0.06 $\rightarrow$ -0.03). On singing voice domain, we observe that using only singing voice data for pre-training yields lower intelligibility compared to using speech data alone (WER: 25.38 vs. 19.39), likely due to the smaller size of the singing voice corpus. Furthermore, when combining speech and singing voice data for pre-training, Vevo2 demonstrates superior intelligibility (WER: 19.39 $\rightarrow$ 15.78) and improved naturalness and similarity (N-CMOS: -1.15 $\rightarrow$ -0.66, SS-CMOS: -0.65 $\rightarrow$ -0.17). These findings provide strong evidence that speech and singing voice generation can be mutually beneficial through unified pre-training. (2) Regarding the effect of post-training on Vevo2, our proposed approach brings further overall improvements to the pre-trained model, including enhanced intelligibility, speaker similarity, and naturalness. This demonstrates the effectiveness and significant potential of post-training techniques in controllable speech and singing voice generation. (3) Compared to existing baselines, the post-trained Vevo2 matches and even outperforms them on certain metrics in the expressive speech domain, despite its lower frame rate in the content-style tokenizer. More significantly, on singing voice domain, Vevo2 substantially outperforms existing zero-shot TTS models in subjective evaluations (N-CMOS: gaps greater than 1.5; SS-CMOS: gaps greater than 0.9), benefiting greatly from the incorporation of singing voice data during training. This validates Vevo2’s effectiveness as a unified controllable speech and singing voice generation model.
4.3 Controllability of Vevo2
Table 2: Results on zero-shot SVS, speech editing, and singing lyric editing tasks.
Zero-Shot Singing Voice Synthesis
Model English Chinese
WER SIM
N-CMOS
Melody-MOS
WER SIM
N-CMOS
Melody-MOS
Ground Truth 17.89 0.640 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 7.40 0.675 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
StyleSinger Zhang et al. (2024d) - - - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 14.04 0.555 -0.37 $_{\scriptscriptstyle\pm\text{0.16}}$ 2.33 $_{\scriptscriptstyle\pm\text{0.36}}$
TCSinger Zhang et al. (2024c) 40.83 0.605 -0.09 $_{\scriptscriptstyle\pm\text{0.12}}$ 2.55 $_{\scriptscriptstyle\pm\text{0.27}}$ 14.26 0.651 -0.28 $_{\scriptscriptstyle\pm\text{0.17}}$ 2.58 $_{\scriptscriptstyle\pm\text{0.19}}$
Vevo2-base 60.83 0.586 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 23.60 0.653 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
Vevo2 24.77 0.643 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 2.29 $_{\scriptscriptstyle\pm\text{0.26}}$ 9.83 0.669 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 2.44 $_{\scriptscriptstyle\pm\text{0.18}}$
Zero-Shot Editing
Model Expressive Speech Singing Voice
WER SIM N-CMOS PS-CMOS WER SIM N-CMOS MS-CMOS
SSR-Speech Wang et al. (2025b) 29.56 0.588 -1.36 $_{\scriptscriptstyle\pm\text{0.09}} -$0.89 $_{\scriptscriptstyle\pm\text{0.14}}$ 51.92 0.710 -1.43 $_{\scriptscriptstyle\pm\text{0.17}} -$1.10 $_{\scriptscriptstyle\pm\text{0.08}}$
F5-TTS Chen et al. (2025) 21.35 0.733 -0.19 $_{\scriptscriptstyle\pm\text{0.12}} -$0.21 $_{\scriptscriptstyle\pm\text{0.16}}$ 29.78 0.784 -1.15 $_{\scriptscriptstyle\pm\text{0.23}} -$0.97 $_{\scriptscriptstyle\pm\text{0.19}}$
Vevo2-base 23.54 0.795 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 29.69 0.841 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
Vevo2 16.83 0.799 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 17.98 0.848 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$
    Vevo2-base is pre-trained on both speech and singing data, corresponding to Vevo2-base (101K, 7K) in Table 1.
Text and Prosody Control To evaluate Vevo2’s controllability of text and prosody (melody), we conduct various tasks including zero-shot SVS, speech editing, singing lyric editing, humming-to-singing, and instrument-to-singing (the latter two tasks will be discussed in Section 4.4).
In the context of zero-shot SVS, given a target text, a music notation like MIDI, and a reference singer’s waveform, the model is required to generate singing voice that matches the content of the target text and music notation while preserving the timbre and style of the reference singer Zhang et al. (2024d, c); Zhao et al. (2025). For Vevo2, we render MIDI into piano waveform and use it as the prosodic source, which can also be considered as instrument-to-sing. Note that we also conduct instrument-to-sing experiments using real instrumental sound recordings, which will be discussed in Section 4.4. The results are shown in Table 2, where we include StyleSinger Zhang et al. (2024d) and TCSinger Zhang et al. (2024c) as two baselines. It demonstrates that: (1) After post-training, Vevo2 shows improved text-following ability and speaker similarity in both English and Chinese. We suspect that this is because that during pre-training, Vevo2 had not encountered prosody tokens extracted from instrumental sounds, resulting in unstable performance. By post-training, we not only supplement with MIDI-rendered instrumental sound data, but also adopt the specific alignment learning, effectively eliciting the model’s potential. (2) Compared to the baselines, the post-trained Vevo2 demonstrates superior text-following ability (WER) and naturalness (N-CMOS). Regarding melody-following, although the post-trained Vevo2’s Melody-MOS is slightly lower than TCSinger’s, its scores for both English and Chinese exceed 2.0, indicating that Vevo2 is at least capable of “roughly following the melody contour” (as defined by Melody-CMOS score of 2.0) in most cases.
In addition to SVS, we adopt editing tasks to evaluate Vevo2’s text and prosody controllability. Vevo2’s advantage in editing tasks stems from its use of prosody tokens as explicit control, enabling it to preserve the prosody (melody) of the raw voice during editing. For instance, in singing lyric editing tasks Zhang et al. (2022c); Hong et al. (2023), Vevo2 can modify only the lyrics while maintaining the original singing voice melody. The results are shown in Table 2, where we include two speech editing baselines, SSR-Speech Wang et al. (2025b) and F5-TTS Chen et al. (2025). It can be observed that compared to the baselines, the post-trained Vevo2 not only achieves better intelligibility (WER), naturalness (N-CMOS), and speaker similarity (SIM), but also demonstrates superior prosody preservation of the original voice (PS-CMOS), particularly in the singing voice domain (MS-CMOS). We will provide more audio samples at our website to further demonstrate Vevo2’s advantages in singing lyric editing tasks.
Table 3: Results on zero-shot SVC task.
Model English Chinese
WER SIM
N-
CMOS
Style-
CMOS
Melody-
MOS
WER SIM
N-
CMOS
Style-
CMOS
Melody-
MOS
Ground Truth 16.48 - - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 12.82 - - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
FACodec Ju et al. (2024) 32.81 0.434 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 36.97 0.473 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
Vevo-FM Zhang et al. (2025a) 24.05 0.567 -0.64 $_{\scriptscriptstyle\pm\text{0.07}} -$0.45 $_{\scriptscriptstyle\pm\text{0.16}}$ 0.71 $_{\scriptscriptstyle\pm\text{0.21}}$ 22.85 0.610 -0.55 $_{\scriptscriptstyle\pm\text{0.19}} -$0.44 $_{\scriptscriptstyle\pm\text{0.08}}$ 0.88 $_{\scriptscriptstyle\pm\text{0.27}}$
CosyVoice2-FM Du et al. (2024b) 23.59 0.553 -0.64 $_{\scriptscriptstyle\pm\text{0.11}} -$0.63 $_{\scriptscriptstyle\pm\text{0.18}}$ 2.25 $_{\scriptscriptstyle\pm\text{0.15}}$ 20.14 0.589 -0.63 $_{\scriptscriptstyle\pm\text{0.15}} -$0.54 $_{\scriptscriptstyle\pm\text{0.07}}$ 1.65 $_{\scriptscriptstyle\pm\text{0.23}}$
NeuCoSVC 2 Sha et al. (2024) 30.61 0.481 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 32.26 0.519 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
SeedVC Liu (2024) 22.86 0.508 -0.05 $_{\scriptscriptstyle\pm\text{0.13}} -$0.24 $_{\scriptscriptstyle\pm\text{0.15}}$ 2.89 $_{\scriptscriptstyle\pm\text{0.09}}$ 15.65 0.550 -0.33 $_{\scriptscriptstyle\pm\text{0.21}} -$0.54 $_{\scriptscriptstyle\pm\text{0.14}}$ 2.93 $_{\scriptscriptstyle\pm\text{0.14}}$
Vevo2-FM 29.82 0.587 -0.12 $_{\scriptscriptstyle\pm\text{0.16}} -$0.11 $_{\scriptscriptstyle\pm\text{0.12}}$ 2.91 $_{\scriptscriptstyle\pm\text{0.12}}$ 22.54 0.611 -0.03 $_{\scriptscriptstyle\pm\text{0.30}} -$0.32 $_{\scriptscriptstyle\pm\text{0.17}}$ 2.90 $_{\scriptscriptstyle\pm\text{0.08}}$
Vevo2 11.64 0.601 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 2.24 $_{\scriptscriptstyle\pm\text{0.18}}$ 14.53 0.623 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 2.38 $_{\scriptscriptstyle\pm\text{0.25}}$
    Vevo2-FM denotes the flow-matching model of Vevo2. The best and the second best result (excluding ground truth) is shown in bold and by underlined.
Timbre and Style Control To evaluate Vevo2’s controllability of timbre and style, we conduct various tasks including zero-shot voice conversion (VC), singing voice conversion (SVC), speech style conversion (for accent and emotion), and singing style conversion. The results of zero-shot SVC task are presented in Table 3, with other experimental results provided in Appendix C.
For zero-shot SVC, we include five baselines: three models pre-trained only on speech data (FACodec Ju et al. (2024), Vevo Zhang et al. (2025a), and CosyVoice 2 Du et al. (2024b)), and two models pre-trained on both speech and singing voice data (NeuCoSVC 2 Sha et al. (2024) and SeedVC Liu (2024)). Notably, for Vevo and CosyVoice 2, we utilize their flow-matching (FM) models. Similarly, we include Vevo2’s FM model as a baseline (denoted as Vevo2-FM). The results in Table 3 reveal two key findings: (1) Comparing Vevo2-FM with the existing baselines, Vevo2-FM demonstrates superior timbre-disentanglement (higher SIM for both English and Chinese), validating its effectiveness for timbre control. Additionally, it exhibits strong melody modeling capability (high Melody-MOS), attributable to the explicit modeling of chromagram features in our content-style tokenizer. However, due to the lower frame rate of content-style tokens (12.5 Hz), Vevo2-FM shows slightly lower intelligibility compared to Vevo-FM (50 Hz), CosyVoice2-FM (25 Hz), and SeedVC (using continuous features). (2) Comparing Vevo2-FM and Vevo2, there are two interesting observations: a) Vevo2 shows significant improvements in intelligibility, which we attribute to both the explicit use of text input and intelligibility alignment during post-training. b) Although Vevo2’s Melody-MOS is lower than Vevo2-FM’s, its absolute value still exceeds 2, indicating its ability to “roughly follow the melody contour”. This trade-off benefits style-converted SVC [^3] For style-converted SVC, we aim for the generated singing voice to follow the source’s coarse-grained melody contour while imitating fine-grained expression details (such as vibrato) from the reference Zhang et al. (2025a); Violeta et al. (2025).: specifically, Vevo2 achieves better style imitation performance (higher Style-CMOS), leading to improved speaker similarity (higher SIM). Similar characteristics are also observed in style-converted VC tasks Zhang et al. (2025a).
4.4 Effectiveness of Multi-Objective Post-Training
To explore the individual effects of intelligibility alignment and prosody similarity alignment, we conduct an ablation study on Vevo2’s post-training process. Specifically, based on the pre-trained model (i.e., Vevo2-base), we introduce a post-trained model that exclusively uses intelligibility reward to optimize, denoted as Base w/ Intell. We use two challenging tasks to evaluate: humming-to-singing and instrument-to-singing, which also demonstrate Vevo2’s unique capabilities in prosody control. For the instrument-to-sing task, we design two subjective metrics (see Appendix D for details): given a target text, a target melody (i.e., the instrument sound), and a generated voice, subjects perform binary classification to determine whether the voice accurately sings the text (text accuracy) and the melody (melody accuracy). The results shown in Figure 3 reveal several key findings. When using only the intelligibility reward, the model demonstrates significant improvement in intelligibility (WER and Text Accuracy). However, this single-objective optimization potentially compromises other objectives, as evidenced by Base w/ Intell’s Melody Accuracy dropping even below that of the pre-trained model (from 65.0% to 50.0%). Interestingly, when employing both rewards jointly, we observe not only the enhanced melody-following capability but also the improved text-following ability (e.g., text accuracy increases from 75.0% to 90.0%). We hypothesize that this improvement stems from strengthened prosody modeling, which in turn benefits pronunciation and consequently enhances intelligibility. This further verifies the advantages of our proposed multi-objective post-training approach.
Model Humming-to-Singing Instrument-to-Singing
WER SIM FPC WER SIM FPC
Base 32.86 0.534 0.769 40.38 0.503 0.716
       w/ Intell 17.49 0.581 0.774 20.03 0.502 0.731
          w/ Prosody 17.27 0.585 0.784 17.94 0.525 0.745
    The first and last model represent Vevo2-base and Vevo2, respectively. w/ Intell: Only intelligibility reward is used for post-training.
Figure 3: Effect of intelligibility reward (w/ Intell) and prosody similarity reward (w/ Prosody) for post-training. The right figure presents subjective evaluation results on the instrument-to-sing task.
5 Conclusion
This paper introduces Vevo2, a framework that bridges controllable speech and singing voice generation through unified prosody learning. Our approach leverages two proposed tokenizers—a notation-free prosody tokenizer that captures prosody and melody from diverse audio sources, and a low-frame-rate and timbre-disentangled content-style tokenizer—along with unified pre-training strategies and a multi-objective post-training phase. Vevo2 demonstrates effectiveness and versatile controllability, achieving strong generalization across a wide array of synthesis, conversion, and editing tasks for both speech and singing, and uniquely enabling novel applications like humming-to-singing and instrument-to-singing. These findings verify the mutual benefits of unified modeling for both speech and singing voice, highlighting promising directions for future research in further unifying diverse audio domains and advancing multi-faceted controllability.
Appendix A Details of Vevo2
A.1 Prosody Tokenizer and Content-Style Tokenizer
This paper processes and trains on speech and singing voice data with a sampling rate of 24 kHz. For chromagram feature extraction, we employ Librosa’s implementation [^4] https://librosa.org/doc/main/generated/librosa.feature.chroma_stft.html with the following parameters: number of FFT points set to 1920, hop length of 480, window size of 1920, and 24 chroma bins. For Whisper hidden features Radford et al. [2023], we utilize the encoder output from the pre-trained whisper-medium model [^5] https://huggingface.co/openai/whisper-medium. We adopt the VQ-VAE implementations from Amphion [^6] https://github.com/open-mmlab/Amphion/tree/main/models/vc/vevo Zhang et al. [2024a], Li et al. [2025], which serves for MaskGCT Wang et al. [2025a] and Vevo Zhang et al. [2025a]. Our only modification is the addition of a CNN-based downsampling layer before the encoder input and a CNN-based upsampling layer after the decoder output. The extracted chromagram features and Whisper features are both at 50 Hz. For the prosody tokenizer, the down(up)-sampling ratio is 8, resulting in a token rate of 6.25 Hz, with a codebook vocabulary size of 512. For the content-style tokenizer, the down(up)-sampling ratio is 4, yielding a token rate of 12.5 Hz, with a codebook vocabulary size of 16,384.
All models in this study use AdamW Kingma and Ba [2015], Loshchilov and Hutter [2019] as the optimizer. For the two tokenizers, we use a peak learning rate of 1e-4, with linear warm-up for 10K steps followed by decay over the remaining training period. Both tokenizers are trained for 300K updates. The prosody tokenizer and content-style tokenizer contain 38M and 44M parameters, respectively.
A.2 Content-Style Modeling
During the content-style modeling stage, we initialize the AR model using Qwen2.5-0.5B [^7] https://huggingface.co/Qwen/Qwen2.5-0.5B Yang et al. [2024b] and expand its original vocabulary to accommodate our proposed prosody tokens and content-style tokens. The total parameter count of the AR model is 509M. The text instruction prefixes for IPL ($\bm{I}_{ipl}$) and EPL ($\bm{I}_{epl}$) are defined as: "User will provide you with a text. Please vocalize it with natural expression." and "User will provide you with a text. Please first generate a good prosodic instruction, then vocalize the text based on it." respectively. For the BPE-based text tokenizer, we adopted the same tokenization strategy as CosyVoice 2, which masks out one-to-many tokens to prevent excessive pronunciation length and mitigate corner cases caused by data sparsity Du et al. [2024b]. During pre-training, we use a peak learning rate of 5e-4, with linear warm-up for 32K steps followed by decay over the remaining training period. The model is pre-trained for 500K updates.
A.3 Acoustic Modeling
During the acoustic modeling stage, we adopted the flow-matching (FM) transformer architecture from Vevo. The transformer consists of 16 layers with 16 attention heads and a hidden dimension of 1024. The FM model contains 363M parameters in total. To enhance training efficiency, we incorporated the REPA strategy Yu et al. [2025] and enforced alignment between the hidden output of the transformer’s 5th layer and the W2v-BERT 2.0 features [^8] https://huggingface.co/facebook/w2v-bert-2.0 Chung et al. [2021], Wang et al. [2025a] during the training process. We use a peak learning rate of 7.5e-5, with linear warm-up for 32K steps followed by decay over the remaining training period. The FM model is trained for 700K updates.
To develop a unified vocoder for both speech and singing voice, we fine-tuned Vevo’s speech vocoder [^9] https://huggingface.co/amphion/Vevo Zhang et al. [2025a] using our training data (comprising 101K hours of speech and 7K hours of singing voice data). The vocoder, which is based on the Vocos architecture Siuzdak [2024], contains 255M parameters and underwent fine-tuning for 572K updates.
A.4 Post-training
(a) Pre-training with EPL and IPL.
(b) Reward model training with EPL and IPL.
Figure 4: Explicit Prosody Learning (EPL) and Implicit Prosody Learning (IPL).
A.4.1 Intelligibility Reward Model
For training an intelligibility reward model, we employed the widely-adopted Bradley-Terry (BT) approach Bradley and Terry [1952], Ouyang et al. [2022], Bai et al. [2022]. As illustrated in Figure 4, we utilized a preference corpus (INTP Zhang et al. [2025b] in our study) to train the pairwise reward model using positive and negative samples (Figure 4(b)). Analogous to our pre-training strategy that incorporated both EPL and IPL strategies (Figure 4(a)), we randomly alternated between IPL and EPL with equal probability during reward model training.
The reward model was initialized using our pre-trained AR model. We implemented a learning rate schedule with a peak value of 1e-4, incorporating a linear warm-up phase for the first 5K steps, followed by a decay phase throughout the remaining training period. The training process continued for 55K updates. During inference, the model accepts either a $\bm{T},\bm{Q}_{cs}]$ or $\bm{T},\bm{Q}{p},\bm{Q}_{cs}]$ sequence as input and produces a float value representing the intelligibility reward.
A.4.2 Prosody Similarity Reward Model
For the prosody similarity alignment, we prepared prompt data incorporating both text and prosody tokens. Specifically, we employed three types of prompts: (1) For speech data, we utilized 20K positive samples randomly selected from INTP, retaining their original text and extracting prosody tokens from their corresponding waveforms. (2) For singing voice data, we selected 20K samples from M4Singer Zhang et al. [2022a], similarly preserving their original text and extracting prosody tokens from the original waveforms. (3) For instrumental sounds, we employed 16 MIDI instruments (Table 4).
Table 4: The 16 MIDI instruments used for post-training.
Category Instrument Name
Piano Acoustic Grand Piano
Bright Acoustic Piano
Electric Piano
Strings Violin
Cello
Orchestral Harp
Woodwinds Alto Sax
Flute
Clarinet
Brass Trumpet
Trombone
Folk Pan Flute
Koto
Voice Choir Aahs
Voice Oohs
Synth Voice
During each training iteration with the 20K M4Singer samples, we randomly selected one instrument and rendered the corresponding audio using pretty_midi [^10] https://github.com/craffel/pretty-midi. We use the original text of the sample and extract prosody tokens from these rendered instrumental sounds.
A.5 Inference-Time Control
A.5.1 Duration Control
As discussed in Section 3.5, we discovered that Vevo2 enables effective duration control of generated outputs through the utilization of prosody tokens as input. This capability arises from a fundamental characteristic: both our proposed prosody tokenizer and content-style tokenizer operate at fixed frame rates (6.25 Hz and 12.5 Hz, respectively). During training, this results in a consistent linear relationship between the sequence lengths of content-style tokens and prosody tokens, specifically maintaining a 2:1 ratio. The AR model readily learns this straightforward length pattern. Consequently, during inference, we can control the output duration by applying linear scaling transformations to the chromagram features of the prosodic source—leveraging the inherent robustness of chromagram features to such transformations—thereby modulating the length of prosody tokens.
Table 5: Effectiveness of duration control of Vevo2 for different tasks.
Task DDUR ($\downarrow$) Duration Consistency ($\uparrow$)
Speech Editing 0.22s 97.1%
Singing Lyric Editing 0.20s 98.4%
Singing Voice Conversion (style-converted) 0.12s 98.4%
In our experiments, we investigated Vevo2’s duration control effectiveness across three tasks: speech editing, singing lyric editing, and singing voice conversion. We adopt the DDUR metric Zhang et al. [2020, 2025a] ($\downarrow$), which measures the average absolute difference in seconds between the actual generated duration ($\hat{d}$) and the target duration ($d$), defined as $\text{abs}(\hat{d}-{d})$. Additionally, we introduce duration consistency ($\uparrow$), defined as $\frac{\text{abs}(\hat{d}-{d})}{d}$, to evaluate the relative accuracy of duration control. For editing tasks, following F5-TTS Chen et al. [2025], we calculate the target duration based on the linear relationship between the lengths of edited and raw text. For singing voice conversion, we constrain the target duration to match the source waveform duration. As demonstrated in Table 5, Vevo2 achieved over 97% duration consistency across all three tasks, exhibiting good duration controllability despite its autoregressive nature.
A.5.2 Pitch Region Control
The pitch region control is particularly effective for voice conversion (VC) and singing voice conversion (SVC) tasks Zhang et al. [2024e], Ning et al. [2023], Huang et al. [2023]. For Vevo2, to achieve the pitch region control:
  • In the style-preserved VC or SVC tasks, where only the FM model (Vevo2-FM) is utilized for conversion as shown in Figure 1: $\boldsymbol{\bm{cs}}$, $\boldsymbol{\bm{mel}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{mel}}}$], we can shift the F0 of the source waveform to match the pitch region of the timbre reference before extracting chromagram features. This process yields pitch-region-controlled $\boldsymbol{\bm{cs}}$ (while the whisper features for extraction can remain from the original source waveform).
  • In the style-converted SVC task, as shown in Figure 1: $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$], we similarly shift the F0 of the source waveform to align with the timbre reference’s pitch region before extracting $\boldsymbol{\bm{p}}$.
Table 6: Effect of pitch shift of Vevo2 in VC and SVC tasks.
Model English Chinese
WER SIM FPC WER SIM FPC
Voice Conversion
Vevo2-FM 9.35 0.645 0.635 6.88 0.725 0.746
     w/o pitch shift 9.07 0.607 0.640 6.60 0.693 0.754
Singing Voice Conversion
Vevo2-FM 29.82 0.587 0.826 22.54 0.611 0.878
     w/o pitch shift 27.79 0.538 0.802 20.58 0.545 0.876
Vevo2 11.64 0.601 0.782 14.53 0.623 0.846
     w/o pitch shift 11.67 0.556 0.750 13.15 0.563 0.817
The effectiveness of Vevo2’s pitch shift control in VC and SVC tasks is demonstrated in Table 6. The results reveal two key findings: First, using pitch shift significantly enhances speaker similarity (SIM) for both Vevo2-FM (using only the FM model) and the complete Vevo2 system (utilizing both AR and FM models). Second, we observe a slight increase in WER when applying pitch shift. We hypothesize that this minor degradation occurs because the content-style tokens and prosody tokens extracted from inference-time pitch-shifted audio represent patterns less frequently encountered in the training data. In future work, we plan to incorporate such pitch shift perturbations during training to mitigate these small discrepancies arising from the train-inference mismatch.
Appendix B Details of Baselines
Zero-Shot Text to Speech
  • F5-TTS Chen et al. [2025]: It adopts a single-stage flow-matching transformer to predict Mel spectrograms from text. It is pre-trained on Emilia He et al. [2024] (101K hours of speech data). We use the officially released checkpoint [^11] https://huggingface.co/SWivid/F5-TTS to generate samples.
  • MaskGCT Wang et al. [2025a]: It adopts a two-stage pipeline: text to content-style tokens (50 Hz), and content-style tokens to acoustic tokens. It is pre-trained on Emilia He et al. [2024] (101K hours of speech data). We use the officially released checkpoint [^12] https://huggingface.co/amphion/MaskGCT to generate samples.
  • CosyVoice 2 Du et al. [2024a]: It contains an AR transformer to predict the content-style tokens (25 Hz) from text, and a flow-matching transformer to predict Mel spectrograms. We use the officially released checkpoint [^13] https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B to generate samples, which is pre-trained on 167K hours of private data.
Zero-Shot Singing Voice Synthesis
  • StyleSinger Zhang et al. [2024d]: It is a zero-shot SVS model for style transfer that is pre-trained on the Chinese GTSinger data Zhang et al. [2024b]. We use the officially released checkpoint [^14] https://github.com/AaronZ345/StyleSinger to generate samples.
  • TCSinger Zhang et al. [2024c]: It is a zero-shot SVS model for style transfer across cross-lingual speech and singing styles, along with multi-level style control. It is pre-trained on the GTSinger data Zhang et al. [2024b]. We use the officially released checkpoint [^15] https://github.com/AaronZ345/TCSinger to generate samples.
Zero-Shot Editing
Zero-Shot Voice Conversion and Singing Voice Conversion
  • FACodec Ju et al. [2024]: It adopts an auto-encoder and residual vector quantization based architecture. It decouples the raw waveform into factorized attributes through ASR, F0 prediction, and speaker classification tasks, trained on the Libri-light dataset Kahn et al. [2020]. We use the released checkpoint in Amphion  [^19] https://huggingface.co/amphion/naturalspeech3_facodec Zhang et al. [2024a], Li et al. [2025] (which is implemented by the authors) to generate samples.
  • Vevo-FM Zhang et al. [2025a]: It is the flow-matching model of Vevo (i.e., Vevo-Timbre in the original paper), which uses a 50 Hz VQ-VAE content-style tokenizer based on HuBERT Hsu et al. [2021] representations. It is pre-trained on Emilia (101K hours of speech data). We use the officially released checkpoint [^20] https://huggingface.co/amphion/Vevo to generate samples.
  • CosyVoice2-FM Du et al. [2024b]: It is the flow-matching model of CosyVoice 2 that uses 25 Hz content-style tokenizer, consistent with the configuration described in the zero-shot TTS task above.
  • NeuCoSVC 2 Sha et al. [2024]: It is an enhanced version of NeuCoSVC and is pre-trained on an extensive internal dataset comprising approximately 500 hours of singing voice data and various open-source speech datasets Sha et al. [2024]. We use the officially released checkpoint [^21] https://github.com/thuhcsi/NeuCoSVC/tree/NeuCoSVC2 to generate samples.
  • SeedVC Liu [2024]: It is a diffusion transformer that uses an external timbre shifter during training to perturb the source speech timbre. We use the officially released pre-trained VC and SVC models [^22] https://github.com/Plachtaa/seed-vc to conduct the inference.
Zero-Shot Emotion Conversion and Accent Conversion
  • Vevo Zhang et al. [2025a]: It is named as Vevo-Style in the original paper, which contains an AR transformer to predict content-style tokens (50 Hz) from content tokens, and a flow-matching transformer to predict Mel spectrograms. We use the officially released checkpoint [^23] https://huggingface.co/amphion/Vevo to generate samples, which is pre-trained on 101K hours of Emilia.
Appendix C Additional Experimental Details and Results
C.1 Zero-Shot Text to Speech
The zero-shot TTS evaluation results for expressive speech and singing voice domains are presented in Table 1. Table 7 demonstrates the performance on the regular speech domain using the SeedTTS evaluation sets Anastassiou et al. [2024a], a widely adopted benchmark in the TTS field Wang et al. [2025a], Du et al. [2024b], Chen et al. [2025]. The results corroborate our findings from Section 4.2. Specifically, we observe that incorporating singing voice data alongside speech data enhances performance across all metrics (WER, SIM, UTMOS) in the regular speech domain. Notably, the post-trained Vevo2 achieves comparable performance to existing state-of-the-art baselines, despite utilizing a content-style tokenizer with a lower frame rate.
Table 7: Results of zero-shot TTS task on regular speech domain. The results on expressive speech and singing voice domains are shown in Table 1. (PT?: Whether post-training is applied.)
Model
#Hz
Data
(#hours)
PT?
Regular Speech
WER SIM UTMOS
Ground Truth - - - 1.699 0.740 3.158
F5-TTS Chen et al. [2025] - 101K 3.442 0.670 3.090
MaskGCT Wang et al. [2025a] 50 101K 2.344 0.738 3.136
CosyVoice2 Du et al. [2024a] 25 167K 2.091 0.709 3.685
Vevo2-base
12.5
101K 6.479 0.719 3.378
7K 29.572 0.654 3.081
101K, 7K 5.201 0.720 3.386
Vevo2 12.5 101K, 7K 3.292 0.724 3.405
    #Hz: Frame rate of the content-style tokenizer. Data: Pre-training data, where blue and red denote speech and singing data. The best and the second best result (excluding ground truth) is shown in bold and by underlined.
C.2 Zero-Shot Editing
The zero-shot editing evaluation (Table 2) was conducted using evaluation sets constructed with DeepSeek-V3 DeepSeek-AI et al. [2024]. Specifically, we utilized the same expressive speech and singing voice evaluation data from our zero-shot TTS tasks. For each sample’s text, we generated edited versions using DeepSeek-V3 with the following prompt: "Please edit the following text by adding, deleting, or modifying content as needed. Note: (1) Ignore any existing punctuation errors in the original text and ensure proper punctuation in the edited version. (2) The edited version must differ from the original text. (3) Return only the final edited text without any additional explanations."
C.3 Zero-Shot Voice Conversion
Table 8 presents the results of our voice conversion evaluation, utilizing the same baselines [^24] For SeedVC, we employed the official SVC checkpoint for the zero-shot SVC task (Table 3) and the official VC checkpoint for the zero-shot VC task (Table 8). as the SVC task in Table 3, where we use the SeedTTS VC evaluation set Anastassiou et al. [2024a]. The VC task results lead to conclusions similar to those observed in the SVC experiments: First, Vevo2-FM demonstrates robust timbre disentanglement capabilities, with speaker similarity (SIM) metrics matching or exceeding existing VC models, particularly for English content. Furthermore, the transition from Vevo2-FM to Vevo2 (i.e., from style-preserved VC to style-converted VC) yields additional improvements in speaker similarity. Moreover, the explicit incorporation of text input significantly enhances intelligibility performance.
Table 8: Results on zero-shot VC task.
Model English Chinese
WER SIM UTMOS WER SIM UTMOS
Ground Truth 2.15 - 3.52 1.25 - 2.79
FACodec Ju et al. [2024] 4.88 0.355 2.79 8.20 0.495 2.10
Vevo-FM Zhang et al. [2025a] 3.01 0.536 3.51 4.06 0.695 2.86
CosyVoice2-FM Du et al. [2024b] 4.66 0.530 3.86 2.80 0.728 3.08
NeuCoSVC 2 Sha et al. [2024] 3.57 0.227 3.15 2.90 0.468 2.51
SeedVC Liu [2024] 2.97 0.565 3.31 2.45 0.737 2.69
Vevo2-FM 9.35 0.645 3.59 6.88 0.725 2.83
Vevo2 3.53 0.692 3.81 3.01 0.755 3.00
    Vevo2-FM denotes the flow-matching model of Vevo2. The best and the second best results (excluding ground truth) are shown in bold and underlined.
C.4 Zero-Shot Speech Style Conversion
Table 9 demonstrates Vevo2’s performance on zero-shot accent and emotion conversion tasks. We benchmarked against Vevo Zhang et al. [2025a] and adopted their evaluation sets and two style conversion metrics: Accent-SIM (computed using a pre-trained accent classifier [^25] https://huggingface.co/Jzuluaga/accent-id-commonaccent_ecapa) and Emotion-SIM (calculated via a pre-trained emotion classifier [^26] https://github.com/ddlBoJack/emotion2vec) to assess style conversion performance. The results indicate that Vevo2 achieves comparable style similarity scores to Vevo while exhibiting superior performance in both intelligibility and speaker similarity metrics.
Table 9: Results on zero-shot accent and emotion conversion tasks.
Model Accent Conversion Emotion Conversion
WER SIM Accent-SIM UTMOS WER SIM Emotion-SIM UTMOS
Vevo Zhang et al. [2025a] 30.37 0.702 0.526 3.49 13.31 0.697 0.406 3.95
Vevo2-base 10.08 0.760 0.537 3.72 11.36 0.700 0.400 3.82
Vevo2 7.73 0.768 0.529 3.76 7.89 0.714 0.402 3.89
C.5 Humming-to-Singing and Instrument-to-Singing
The evaluation sets for humming-to-singing and instrument-to-singing tasks in Section 4.4 were constructed through the following procedure. For melody sources, we selected humming samples Liu et al. [2024] and instrumental sounds, including flute Braga Brum [2018] and violin Dong et al. [2022] recordings. For each melody sample, we leveraged its MIDI annotation’s note count to generate appropriate lyrics using DeepSeek-v3. The following prompt was used for DeepSeek-v3: "I will provide the number of notes in a melody. Please write suitable lyrics for it. Note: (1) The number of words in the lyrics should roughly match the number of notes; (2) Please only return the result, without any further explanation." After obtaining both melody and lyrics, we randomly assigned a target singer from SingStyle111 Dai et al. [2023] for each sample, thereby creating evaluation triplets in the form of (melody, lyric, singer). The evaluation set comprises 1,000 samples in total, evenly distributed between English (500) and Chinese (500) lyrics.
Appendix D Subjective Evaluation
We recruited multiple subjects on a compensated basis to conduct subjective evaluations. All participants possessed extensive musical backgrounds, with a minimum of five years of experience in instrumental performance and music production. Each audio sample in our evaluation received at least three independent ratings. We have developed an automated subjective evaluation interface. For each item to be evaluated, users will see three components: the System Interface (i.e., the audio to be evaluated), the Questionnaire, and the Evaluation Criteria.
Naturalness (N-CMOS)
  • System Interface: Users listen to two speech samples, A and B, to compare their naturalness.
  • Questionnaire: Which speech sample sounds more natural and human-like, rather than synthetic or AI-generated?
  • Evaluation Criteria: Options include A +2 (Sample A is much more natural), A +1 (Sample A is slightly more natural), Tie (Both are equally natural), B +1 (Sample B is slightly more natural), and B +2 (Sample B is much more natural).
Speaker Similarity (SS-CMOS)
  • System Interface: Users listen to two speech samples, A and B, to evaluate their similarity to the speech of the reference speaker.
  • Questionnaire: Which speech sounds more like the reference speaker’s voice?
  • Evaluation Criteria: Options include A +2 (Sample A is much more similar), A +1 (Sample A is slightly more similar), Tie (Both are equally similar), B +1 (Sample B is slightly more similar), and B +2 (Sample B is much more similar).
Prosody Similarity (PS-CMOS)
  • System Interface: Users listen to two speech samples, A and B, to evaluate their similarity to the speech of the reference prosody.
  • Questionnaire: Ignore the vocal characteristics (who is speaking), just pay attention to the speaking style (how to speak). Which speech’s speaking style (pace, tone, stress, intonation, pitch, emotion) is more consistent and identical with the reference voice?
  • Evaluation Criteria: Options include A +2 (Sample A is much more similar), A +1 (Sample A is slightly more similar), Tie (Both are equally similar), B +1 (Sample B is slightly more similar), and B +2 (Sample B is much more similar).
Melody Similarity (MS-CMOS)
  • System Interface: Users listen to two singing voice samples, A and B, to evaluate their similarity to the singing voice of the reference melody.
  • Questionnaire: Which singing voice’s melody is more consistent and identical with the reference voice?
  • Evaluation Criteria: Options include A +2 (Sample A is much more similar), A +1 (Sample A is slightly more similar), Tie (Both are equally similar), B +1 (Sample B is slightly more similar), and B +2 (Sample B is much more similar).
Style Similarity (Style-CMOS)
  • System Interface: Users listen to two singing voice samples, A and B, to evaluate their similarity to the singing voice of the reference style.
  • Questionnaire: Which singing voice’s vocal techniques and style (such as the vibrato, falsetto, glissando, etc.) is more similar to the reference voice?
  • Evaluation Criteria: Options include A +2 (Sample A is much more similar), A +1 (Sample A is slightly more similar), Tie (Both are equally similar), B +1 (Sample B is slightly more similar), and B +2 (Sample B is much more similar).
Melody-MOS
  • System Interface: One reference audio, One audio to be evaluated.
  • Questionnaire: Does the singing voice’s melody follow the reference audio?
  • Evaluation Criteria: Options include 1 (Unable to follow), 2 (Roughly following the melody contour), and 3 (Completely following all melody details).
Text Accuracy
  • System Interface: Users listen to the singing voice and compare it to the provided target text to assess whether the voice matches the text.
  • Questionnaire: Does the voice contain any pronunciation errors in the lyrics (such as insertion, omission, or mispronunciation)?
  • Evaluation Criteria: The evaluation is binary: “No Error” (the voice matches the text) or “Has Error” (the voice does not match the text).
Melody Accuracy
  • System Interface: Users listen to the singing voice and compare it to the provided target humming or instrumental melody to assess whether the voice matches the melody.
  • Questionnaire: Does the voice follow the target melody without any out-of-tone errors?
  • Evaluation Criteria: The evaluation is binary: “No Error” (the voice matches the melody) or “Has Error” (the voice does not match the melody).
Appendix E Limitations
While Vevo2 exhibits versatile controllability through reference waveforms, it currently lacks support for natural language-based control mechanisms (e.g., textual prompts such as "sing with a joyful tone") for direct style or prosody manipulation. Instead, it relies exclusively on explicit audio references. Furthermore, the framework’s scope is primarily limited to monophonic melody modeling—following the classical formulation of singing voice synthesis—and may face challenges when processing complex polyphonic instrumental inputs. Future research directions include the integration of language-driven control systems and the extension of capabilities to accommodate more sophisticated musical structures.
Appendix F Broader Impacts
As with other powerful new AI innovations, we recognize this technology brings the potential for misuse and unintended harm. We will build a highly effective classifier that can distinguish between authentic speech and audio generated with Vevo2 to mitigate these possible future risks.